Report on NIRSPEC observing nights of 19 and 23 May 2000

Summary Report on the second and third nights of Gemini/ NIRSPEC 2000A time

The final two nights of Gemini/NIRSPEC observing on Keck II were May 19/20 and May 23/24. The weather was clear with seeing on both nights typically 0.5" as viewed through the K filter.

Mechanically NIRSPEC and the telescope functioned very well; slit rotation was available, offset guiding was much more accurate than on April 8. However, the array was prone to "blotches" occurring at random locations and in random frames. Occasionally these occurred in regions of the array where spectra were being obtained, rendering a portion of that spectral frame unusable. In addition, the array suffered/suffers from persistence - i.e. a faint remnant spectrum of a bright object observed previously persisted on the array through many reads and for as long as an hour. On the first night it took some time for us to recognize this problem and then attempt to work around it; some of the data obtained on this night are compromised by this effect. By the second night we were prepared to deal with it. All of those for whom we obtained spectra of faint objects will notice these residual spectra.

We have learned that the overheads in using NIRSPEC are large. On the night of May 23, during which we observed a combination of bright and faint targets with what we think was pretty high efficiency, the total exposure time was about 5 hours, or roughly 50% of the night. We doubt that we can increase this value significantly unless we devote large fractions of the available time to very long exposures; even on a simple program requiring very long exposures efficiency is unlikely to be greater than 75%. Now that we know this we expect to tailor the 2000B proposal form and the rules for proposals to take it into account.

The following are programs that were attemped, along with a brief summary of the results. Within the next week we will be releasing data to the PI's of the proposals.

023x_f (SDSS mystery object - Fan) - spectra obtained in the J and H bands on May 23/24, believed to be good quality.

028c_t (Cygnus A - Tadhunter) - K band spectrum obtained thru nucleus only on May 23/24, believed to be good quality.

036med (J band droppout in HDF-N - Dickinson) - H band spectrum obtained on May 23/24, believed to be good quality.

040g_d (ERO pair - Dalton) - J and K band spectra obtained of one pair on May 20/21, may be partially deleteriously affected by residuals of previous calibration star spectra.

046mab (Disk galaxy scaling relations - Bershady) - H band echelle spectrum obtained of one target on May 20/21.

048j_m (Oxygen abundances in metal poor stars - Melendez) - echelle spectra of four of the six targets obtained on May 20/21; believed to be good quality.

055m_d (Fe abundance at z=4.7 - Dietrich) - J-H, H, and K spectra obtained on May 20/21 - may be compromised by residual spectra of calibration star.

056rih (V404 Cygni - Hynes) K-band spectrum obtained on May 20/21 - believed to be good quality.

058k_s (CNO abundances in GC/IRS22 - Sellgren) - H and K band echelle spectra obtained on May 23/24; exposures times at H somewhat less than requested.

059s_l (Lines in z=0.5-1.0 galaxies - Lilly) - J band spectrum obtained of one target on May 23/24, believed to be good quality.

---- May 26, 2000

[Science Operations home] [NIRSPEC home]

Last update May 26, 2000; Tom Geballe